• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name gilman

    Log entry time 21:55:30 on November 02, 2008

    Entry number 246322

    Followups:

    keyword=Q3 trip

    Just before 9:30 Q3 left trips. We notice that the He level is 54% and the lead flows are slightly above 100%. Brad is here, so after talking with him I page Scot. While we are talking, the helium liquid level gradually recovers up to ~77%. Scot says that cryo basically controls this and there is little he can do. If it becomes a big problem overnight we likely need to get cryo in. They should look at it in the morning in any case. At this point it was stable and on for ~6 hours since the last trip, so it does not seem to be a big problem. Here there is no loss of beam time since we are still working on the Moller.